# Differential calculus on Hopf–Galois extension via the Đurđević braiding

### v1

@@ 3 Differential calculus generated by the Đurđević braiding
$\pi\big(\mathrm{ver}_{u}^{-1}(A\otimes I_{H})\big)\;=\;\mathrm{can}^{-1}(A\otimes I_{H})\;\subset\;A\otimes_{B}A.$
@@
$\mathrm{ver}_{u}=(\mathrm{id}\otimes\mathrm{pr}_{H^{+}})\circ\mathrm{can}\circ\pi$
@@
Remark 3.2.
Throughout, $I_{H}\subset H^{+}$ denotes a fixed right ideal specifying vertical relations. We neither assume a differential calculus on $H$, nor adjoint coaction invariance of $I_{H}$. Such additional conditions arise only in the framework of quantum principal bundles with compatible calculi on $A$ and $H$, which we do not assume here.
@@
Let $B=A^{\mathrm{co}(H)}\subset A$ be a Hopf–Galois extension with structure Hopf algebra $H$, and let $I_{H}\subset H^{+}$ be a right ideal of $H$. Let $\omega:H\to\Omega^{1}_{\mathrm{u}}(A)$ be the connection $1$-form associated to a strong connection on $A$. Then
@@
Let $B=A^{\mathrm{co}(H)}\subset A$ be a Hopf–Galois extension with structure Hopf algebra $H$, and let $\sigma:A\otimes_{B}A\to A\otimes_{B}A$ be the Đurđević braiding. For a subspace $W\subset A\otimes_{B}A$, denote by $\langle W\rangle_{\sigma}$ the smallest $\sigma$-stable subspace of $A\otimes_{B}A$ containing $W$. Explicitly,
$\langle W\rangle_{\sigma}:=\sum_{n\geq 0}\sigma^{n}(W).$
@@
Let $A$ be a principal right HH–comodule algebra and let $I_{H}\subset H^{+}$ be any fixed right ideal of $H$. Then there exists a right HH–covariant first-order differential calculus
@@
on $A$ which is $\sigma$–generated with respect to $I_{H}$, i.e.
$N_{A}^{\mathrm{bal}}=\big\langle\pi(\omega(I_{H}))\big\rangle_{\sigma}\;\subset\;A\otimes_{A^{\mathrm{co}(H)}}A,$
where $\sigma$ denotes the Đurđević braiding and $\omega$ is a connection 11–form associated to a strong connection on $A$.
@@
Since $A$ is a principal right HH–comodule algebra, it admits a strong connection $\ell:H\to A\otimes A$. Let $\omega:H\to\Omega^{1}_{\mathrm{u}}(A)\subset A\otimes A$ be the associated universal connection 11–form, defined by
$\omega(h^{+})=\ell(h)-\varepsilon(h)\,1_{A}\otimes 1_{A}.$
Let $\pi:A\otimes A\to A\otimes_{A^{\mathrm{co}(H)}}A$ denote the canonical projection and define
$N_{A}^{\mathrm{bal}}:=\big\langle\pi(\omega(I_{H}))\big\rangle_{\sigma},$
the smallest $\sigma$–stable subspace of $A\otimes_{A^{\operatorname{co}H}}A$ containing $\pi(\omega(I_{H}))$.
By construction, $N_{A}^{\mathrm{bal}}$ is $\sigma$–stable and consists of vertical elements. Fix a right HH–covariant AA–subbimodule $N_{A}\subset\Omega^{1}_{\mathrm{u}}(A)$ whose image under $\pi$ equals $N_{A}^{\mathrm{bal}}$. Such a lift exists since $\pi$ is a surjective AA–bimodule map.
@@
By construction, $\Omega^{1}(A)$ is a right HH–covariant first-order differential calculus on $A$, and its balanced defining relations satisfy
$N_{A}^{\mathrm{bal}}=\big\langle\pi(\omega(I_{H}))\big\rangle_{\sigma}.$
Hence $\Omega^{1}(A)$ is $\sigma$–generated with respect to $I_{H}$. ∎
@@ 4 Obstructions to braiding-generated calculi
Let $B=A^{\mathrm{co}(H)}\subset A$ be a Hopf–Galois extension with structure Hopf algebra $H$, and let
@@
be a $\sigma$-generated first-order differential calculus. Then every balanced defining relation in $N_{A}^{\mathrm{bal}}$ is generated from vertical elements via the action of the Đurđević braiding.
@@
Let $\Omega^{1}(A)$ be $\sigma$-generated with respect to a right ideal $I_{H}\subset H^{+}$. Then
@@
Remark 4.3.
Unlike the bicovariant setting of Woronowicz [15], where first–order differential calculi on a Hopf algebra $H$ are classified by right adjoint coaction invariant ideal $I\subset H^{+}$, $\sigma$–generation does not, in general, force the defining subspace $I_{H}\subset H^{+}$ to be a right adjoint coaction invariant ideal of $H$. This reflects the fact that $\sigma$–generated calculi form a strictly more general class than bicovariant differential calculi, as they require only right HH–covariance and compatibility with the Đurđević braiding rather than full bicovariance.
@@ 5 The quantum Hopf fibration over the standard Podleś sphere
Let $A=\mathcal{O}_{q}(SU(2))$ and $H=\mathcal{O}(U(1))$. For any right ideal $I_{H}\subset H^{+}$, there exists a right HH–covariant first-order differential calculus
@@
By Proposition 5.1, AA is a principal right HH–comodule algebra. Hence AA admits a strong connection and an associated connection 11–form $\omega:H\to\Omega^{1}_{\mathrm{u}}(A)$. The statement now follows directly from the general existence theorem for $\sigma$–generated calculi established in Proposition 3.7. ∎
@@ Acknowledgements
The author would like to thank Thomas Weber for suggesting the reference for Đurđević braiding.

### v2

@@ 3 Differential calculus generated by the Đurđević braiding
$\pi\big(\mathrm{ver}_{\mathrm{u}}^{-1}(A\otimes I_{H})\big)\;=\;\mathrm{can}^{-1}(A\otimes I_{H})\;\subset\;A\otimes_{B}A.$
@@
$\mathrm{ver}_{\mathrm{u}}=(\mathrm{id}\otimes\mathrm{pr}_{H^{+}})\circ\mathrm{can}\circ\pi$
@@
Let $B=A^{\mathrm{co}(H)}\subset A$ be a Hopf–Galois extension with structure Hopf algebra $H$, and let $I_{H}\subset H^{+}$ be a right ideal of $H$. Let $\omega:H^{+}\to\Omega^{1}_{\mathrm{u}}(A)$ be the connection $1$-form associated to a strong connection on $A$. Then
@@
Let $B=A^{\mathrm{co}(H)}\subset A$ be a Hopf–Galois extension with structure Hopf algebra $H$, and let $\sigma:A\otimes_{B}A\to A\otimes_{B}A$ be the Đurđević braiding. For a subspace $W\subset A\otimes_{B}A$, denote by $\langle W\rangle_{\sigma}$ the smallest $\sigma$-stable subspace of $A\otimes_{B}A$ containing $W$.
@@
Lemma 3.6.
Let $A$ be a principal right HH–comodule algebra and let $I_{H}\subset H^{+}$ be a right ideal invariant under the right adjoint coaction of $H$. Let $\omega:H^{+}\to\Omega^{1}_{\mathrm{u}}(A)$ be the connection 11–form associated to a strong connection on $A$. Define
$N_{A}^{\mathrm{bal}}=\langle\pi(\omega(I_{H}))\rangle_{\sigma}\subset A\otimes_{A^{\mathrm{co}H}}A.$
Then $N_{A}^{\mathrm{bal}}$ is invariant under the right HH–coaction, that is,
$\delta(N_{A}^{\mathrm{bal}})\subset N_{A}^{\mathrm{bal}}\otimes H.$
Proof.
By Definition 2.4, the connection 11–form $\omega$ satisfies
$\delta_{\Omega^{1}_{\mathrm{u}}(A)}\circ\omega=(\omega\otimes\mathrm{id})\circ\mathrm{Ad}_{R},$
where $\delta_{\Omega^{1}_{\mathrm{u}}(A)}$ denotes the right HH–coaction on $\Omega^{1}_{\mathrm{u}}(A)$ and $\mathrm{Ad}_{R}$ is the right adjoint coaction of $H$.
Since $I_{H}$ is invariant under $\mathrm{Ad}_{R}$, we have
$\mathrm{Ad}_{R}(I_{H})\subset I_{H}\otimes H.$
Hence,
$\delta_{\Omega^{1}_{\mathrm{u}}(A)}(\omega(I_{H}))=(\omega\otimes\mathrm{id})(\mathrm{Ad}_{R}(I_{H}))\subset\omega(I_{H})\otimes H.$
Therefore, $\omega(I_{H})$ is a right HH–subcomodule of $\Omega^{1}_{\mathrm{u}}(A)$.
Let
$\pi:\Omega^{1}_{\mathrm{u}}(A)\longrightarrow A\otimes_{A^{\mathrm{co}H}}A$
be the canonical projection. Since $\pi$ is right HH–colinear, it follows that
$\delta(\pi(\omega(I_{H})))=(\pi\otimes\mathrm{id})(\delta_{\Omega^{1}_{\mathrm{u}}(A)}(\omega(I_{H})))\subset\pi(\omega(I_{H}))\otimes H.$
Thus, $\pi(\omega(I_{H}))$ is a right HH–subcomodule of $A\otimes_{A^{\mathrm{co}H}}A$.
Moreover, the Durdević braiding $\sigma$ is right HH–colinear. Consequently, for all $x\in A\otimes_{A^{\mathrm{co}H}}A$,
$\delta(\sigma(x))=(\sigma\otimes\mathrm{id})(\delta(x)).$
It follows that the right HH–coaction preserves the $\sigma$–stable subspace generated by $\pi(\omega(I_{H}))$.
Since $N_{A}^{\mathrm{bal}}$ is, by definition, the smallest $\sigma$–stable subspace containing $\pi(\omega(I_{H}))$, we conclude that
$\delta(N_{A}^{\mathrm{bal}})\subset N_{A}^{\mathrm{bal}}\otimes H.$
This completes the proof. ∎
@@
Let $A$ be a principal right HH–comodule algebra and let $I_{H}\subset H^{+}$ be a right ideal invariant under the right adjoint coaction of $H$. Assume that there exists a right HH–colinear map
$s:A\otimes_{A^{\mathrm{co}H}}A\longrightarrow\Omega^{1}_{\mathrm{u}}(A)$
such that $\pi\circ s=\mathrm{id}$, where $\pi:\Omega^{1}_{\mathrm{u}}(A)\to A\otimes_{A^{\mathrm{co}H}}A$ is the canonical projection. Then there exists a right HH–covariant first–order differential calculus
@@
on $A$ which is $\sigma$–generated with respect to $I_{H}$.
@@
Let $N_{A}^{\mathrm{bal}}\subset A\otimes_{A^{\mathrm{co}H}}A$ be defined by
$N_{A}^{\mathrm{bal}}=\langle\pi(\omega(I_{H}))\rangle_{\sigma}.$
By Lemma 3.6, $N_{A}^{\mathrm{bal}}$ is a right HH–subcomodule.
Consider the short exact sequence of right HH–comodules
$0\longrightarrow A\Omega^{1}(B)A\longrightarrow\Omega^{1}_{\mathrm{u}}(A)\xrightarrow{\;\pi\;}A\otimes_{A^{\mathrm{co}H}}A\longrightarrow 0.$
By assumption, this sequence admits a right HH–colinear splitting $s$.
@@
$N_{A}:=s(N_{A}^{\mathrm{bal}})\oplus A\Omega^{1}(B)A\subset\Omega^{1}_{\mathrm{u}}(A).$
Since $s$ is right HH–colinear and $A\Omega^{1}(B)A$ is a right HH–subcomodule, it follows that $N_{A}$ is a right HH–subcomodule of $\Omega^{1}_{\mathrm{u}}(A)$. Moreover, by construction,
$\pi(N_{A})=N_{A}^{\mathrm{bal}}.$
Since $\pi$ is an AA–bimodule map and $N_{A}^{\mathrm{bal}}$ is an AA–subbimodule, the subspace $N_{A}$ is an AA–subbimodule of $\Omega^{1}_{u}(A)$. We define the quotient
@@
Because $N_{A}$ is a right HH–covariant AA–subbimodule, the quotient $\Omega^{1}(A)$ inherits a right HH–coaction and hence is a right HH–covariant first–order differential calculus on $A$.
Finally, by construction,
$N_{A}^{\mathrm{bal}}=\pi(N_{A})=\langle\pi(\omega(I_{H}))\rangle_{\sigma},$
which shows that $\Omega^{1}(A)$ is $\sigma$–generated with respect to $I_{H}$. This completes the proof. ∎
Remark 3.8.
The existence of a right HH–colinear splitting $s:A\otimes_{A^{\mathrm{co}H}}A\to\Omega^{1}_{\mathrm{u}}(A)$ means that the short exact sequence of right HH–comodules
$0\longrightarrow A\Omega^{1}(B)A\longrightarrow\Omega^{1}_{\mathrm{u}}(A)\xrightarrow{\;\pi\;}A\otimes_{A^{\mathrm{co}H}}A\longrightarrow 0$
splits equivariantly. This condition can be interpreted as the existence of an equivariant horizontal lift of universal 11–forms.
Such a splitting exists, for example, if $H$ is cosemisimple, in which case the category of right HH–comodules is semisimple. More generally, it holds in many standard examples of quantum principal bundles.
@@ 4 Obstructions to braiding-generated calculi
Let $A$ be a principal right $H$-comodule algebra, and $B=A^{\mathrm{co}(H)}\subset A$. Let
@@
be a $\sigma$-generated first-order differential calculus with respect to a right ideal $I_{H}\subset H^{+}$. Then every balanced defining relation in $N_{A}^{\mathrm{bal}}$ is generated from vertical elements via the action of the Đurđević braiding.
@@
Let $A$ be a principal right $H$-comodule algebra. We consider $\Omega^{1}(A)$ be $\sigma$-generated with respect to a right ideal $I_{H}\subset H^{+}$. Then
@@ 5 The quantum Hopf fibration over the standard Podleś sphere
Let $A=\mathcal{O}_{q}(SU(2))$ and $H=\mathcal{O}(U(1))=\mathbb{C}[z,z^{-1}]$. We consider right ideal $I_{H}=\langle z-1\rangle\subseteq H^{+}$, there exists a right HH–covariant first-order differential calculus
@@
By Proposition 5.1, $A$ is a principal right HH–comodule algebra. Hence $A$ admits a strong connection and an associated connection 11–form $\omega:H\to\Omega^{1}_{\mathrm{u}}(A)$. Since $H$ is cosemisiple, therefore existence of splitting map $s:A\otimes_{B}A\rightarrow\Omega^{1}_{\mathrm{u}}(A)$ is evident.
Moreover,
$\mathrm{Ad}_{R}(z-1)=(z-1)\otimes z+1\otimes(z-1)\in I_{H}\otimes H$
Hence, the statement now follows directly from the general existence theorem for $\sigma$–generated calculi established in Proposition 3.7. ∎
@@
Description of $\sigma$-generated differential calculus
Consider the quantum Hopf fibration
$A=\mathcal{O}_{q}(SU_{2}),\qquad H=\mathcal{O}(U(1))=\mathbb{C}[z,z^{-1}],\qquad B=\mathcal{O}_{q}(S^{2}),$
and let $I_{H}=\langle z-1\rangle\subseteq H^{+}$. In this case, the associated connection 11–form satisfies
$\omega(z-1)=\ell(z)-1\otimes 1,$
Using the explicit form of $\ell$ for the Hopf fibration, one obtains
$\omega(z-1)=\alpha\otimes\alpha^{*}+\gamma\otimes\gamma^{*}-1\otimes 1,$
where $\alpha,\gamma$ denote the standard generators of $\mathcal{O}_{q}(SU_{2})$.
Applying the canonical projection $\pi:\Omega^{1}_{\mathrm{u}}(A)\to A\otimes_{B}A$, we obtain
$\pi(\omega(z-1))=\alpha\otimes_{B}\alpha^{*}+\gamma\otimes_{B}\gamma^{*}-1\otimes_{B}1.$
Hence the balanced submodule $N_{A}^{\mathrm{bal}}$ is generated, as a $\sigma$–stable AA–subbimodule, by the element
$\alpha\otimes_{B}\alpha^{*}+\gamma\otimes_{B}\gamma^{*}-1\otimes_{B}1.$
In this example, the Durdević braiding $\sigma$ is compatible with the $\mathbb{Z}$–grading induced by the U(1)U(1)–coaction. Consequently, the $\sigma$–closure of the above generator does not introduce additional independent relations, and $N_{A}^{\mathrm{bal}}$ corresponds to the vertical direction of the fibration.
Since $H$ is cosemisimple, therefore, the existence of a right HH–colinear splitting $s:A\otimes_{B}A\to\Omega^{1}_{\mathrm{u}}(A)$ is evident, the subbimodule $N_{A}\subset\Omega^{1}_{\mathrm{u}}(A)$ can be chosen as
$N_{A}=s(N_{A}^{\mathrm{bal}})\oplus A\Omega^{1}(B)A.$
Thus $N_{A}$ decomposes into a vertical component generated by $\omega(z-1)$ and a horizontal component coming from the kernel of the projection onto $\Omega^{1}(B)$.
Since the canonical covariant calculus on the standard Podleś sphere $\mathcal{O}_{q}(S^{2})$ is two–dimensional and the vertical calculus associated to $I_{H}=\langle z-1\rangle$ is one–dimensional, it follows that the resulting $\sigma$–generated calculus on $\mathcal{O}_{q}(SU_{2})$ is three–dimensional, namely,
$\dim\Omega^{1}(A)=3.$
@@ Acknowledgements
The author would like to thank Thomas Weber, Emanuele Latini, Antonio Del Donno and Giovanni Gava for helpful discussions.